# Gradient Dynamics of Attention: How Cross-Entropy Sculpts Bayesian Manifolds

### v1

@@ Abstract.
Transformers empirically perform precise probabilistic reasoning in carefully constructed “Bayesian wind tunnels” and in large-scale language models, yet the mechanisms by which gradient-based learning creates the required internal geometry remain opaque. We provide a complete first-order analysis of how cross-entropy training reshapes attention scores and value vectors in a transformer attention head. Our core result is an advantage-based routing law for attention scores,
@@
coupled with a responsibility-weighted update for values,
@@
where $u_{i}$ is the upstream gradient at position $i$ and $\alpha_{ij}$ are attention weights. These equations induce a positive feedback loop in which routing and content specialize together: queries route more strongly to values that are above-average for their error signal, and those values are pulled toward the queries that use them. We show that this coupled specialization behaves like a two-timescale EM procedure: attention weights implement an E-step (soft responsibilities), while values implement an M-step (responsibility-weighted prototype updates), with queries and keys adjusting the hypothesis frame. Through controlled simulations, including a sticky Markov-chain task where we compare a closed-form EM-style update to standard SGD, we demonstrate that the same gradient dynamics that minimize cross-entropy also sculpt the low-dimensional manifolds identified in our companion work as implementing Bayesian inference. This yields a unified picture in which optimization (gradient flow) gives rise to geometry (Bayesian manifolds), which in turn supports function (in-context probabilistic reasoning).
@@ 1. Introduction
Transformers have become the dominant architecture for sequence modeling, yet we still lack a mechanistic understanding of how gradient descent sculpts their internal representations. Recent work has shown that in controlled “Bayesian wind tunnels” a small transformer can reproduce analytic posteriors exactly, with keys forming orthogonal hypothesis axes, queries implementing progressive belief updates, and values unfurling along one-dimensional manifolds parameterized by posterior entropy (concurrent_bayes). Our companion scaling paper extends this picture to production-scale models, showing similar geometric signatures in Pythia, Phi-2, and LLaMA.
@@
This paper provides that missing link. We analyze a single-head attention block trained with cross-entropy and derive all first-order gradients with respect to scores $s_{ij}$, queries $q_{i}$, keys $k_{j}$, and values $v_{j}$. We then interpret the resulting gradient flows and connect them to the Bayesian manifolds observed empirically.
@@ 1.1. Contributions
Taken together with (concurrent_bayes), our results provide a unified story:
@@ 5.4. Analogy Table
Crucially, however, our companion work (concurrent_bayes) shows that the point-estimate parameters learned in this way support Bayesian computation in representation space: value manifolds, key frames, and query trajectories implement Bayesian belief updates in context. The present paper explains why cross-entropy and gradient descent naturally create these structures.
@@ 6.1. Value Manifold Unfurling
In wind-tunnel experiments (concurrent_bayes), we observe that:
@@ Results.
Figures Figure 7Figure 9 plot loss, accuracy, and entropy as a function of training step. As expected, EM and SGD methods converge to the same solution. This may not always happen due to local minima. EM converges much faster, with loss and entropy dropping sharply and accuracy rising quickly, while SGD progresses more slowly. It is interesting to note that even with this single layer and head architecture, the final empirical predictive entropy is very close to the theoretically minimum Bayesian entropy as described in Table 1. The theoretical entropy is higher due to addition of noise in defining $x_{t+1}$.
@@ 9.1. Bayesian Interpretations of Transformers
Several works argue that transformers implement approximate Bayesian inference, either behaviorally or via probing (e.g. xie2022explanation; von2023transformers). Our companion paper (concurrent_bayes) demonstrates exact Bayesian behavior and geometric signatures in small wind tunnels, while a scaling paper shows similar patterns in production LLMs. The present work explains how gradient dynamics produce these geometries. It further shows that minimum theoretical Bayesian predictive entropy is close to the empirical entropy. This is studied in greater detail in the companion paper.
@@ 9.2. Mechanistic Interpretability
Mechanistic interpretability studies identify specific heads and circuits performing copy, induction, and other algorithmic tasks (elhage2021mathematical; olsson2022context). Our framework complements this line by explaining how specialization arises from the interaction between routing and content, rather than treating specialized heads as primitive.
@@ 9.3. Optimization and Implicit Bias
The implicit bias of gradient descent in linear and deep networks has been extensively studied (soudry2018implicit; ba2022high). We extend these ideas to attention: gradient descent implicitly favors representations where routing aligns with error geometry and values lie on low-dimensional manifolds that support Bayesian updating.

### v2

@@ Abstract.
Transformers empirically perform precise probabilistic reasoning in carefully constructed “Bayesian wind tunnels” and in large-scale language models, yet the mechanisms by which gradient-based learning creates the required internal geometry remain opaque. We provide a systematic first-order analysis of how cross-entropy training reshapes attention scores and value vectors in a transformer attention head. Our core result is an advantage-based routing law for attention scores,
@@
coupled with a responsibility-weighted update for values $v_{j}$,
@@
where $u_{i}$ is the upstream gradient at position $i$ and $\alpha_{ij}$ are attention weights. These equations induce a positive feedback loop in which routing and content specialize together: queries route more strongly to values that are above-average for their error signal, and those values are pulled toward the queries that use them. We show that this coupled specialization behaves like a two-timescale EM procedure: attention weights implement an E-step (soft responsibilities), while values implement an M-step (responsibility-weighted prototype updates), with queries and keys adjusting the hypothesis frame. Through controlled simulations, including a sticky Markov-chain task where we compare a closed-form EM-style update to standard SGD, we demonstrate that the same gradient dynamics that minimize cross-entropy also sculpt the low-dimensional manifolds identified in our companion work as implementing Bayesian inference. Further, in EM versus SGD comparison, EM converges much faster, and predictive entropy for both is close to minimum Bayesian entropy. This yields a unified picture in which optimization (gradient flow) gives rise to geometry (Bayesian manifolds), which in turn supports function (in-context probabilistic reasoning).
@@ 1. Introduction
Transformers have become the dominant architecture for sequence modeling, yet we still lack a mechanistic understanding of how gradient descent sculpts their internal representations. Recent work has shown that in controlled “Bayesian wind tunnels” a small transformer can reproduce analytic posteriors exactly, with keys forming orthogonal hypothesis axes, queries implementing progressive belief updates, and values unfurling along one-dimensional manifolds parameterized by posterior entropy (Agarwal et al., 2025). Our companion scaling paper extends this picture to production-scale models, showing similar geometric signatures in Pythia, Phi-2, and LLaMA.
@@
This paper provides a mechanistic account. We analyze a single-head attention block trained with cross-entropy and derive all first-order gradients with respect to scores $s_{ij}$, queries $q_{i}$, keys $k_{j}$, and values $v_{j}$. We then interpret the resulting gradient flows and connect them to the Bayesian manifolds observed empirically.
@@ 1.1. Contributions
Taken together with (Agarwal et al., 2025), our results provide a unified story:
@@
Clarification on “Bayesian inference.” Throughout this paper, “Bayesian inference” refers to the Bayesian posterior predictive over latent task variables—not a posterior over network weights. We show that cross-entropy training sculpts geometry that implements Bayesian filtering over in-context hypotheses, not Bayesian weight uncertainty.
@@ 5.4. Analogy Table
Crucially, however, our companion work (Agarwal et al., 2025) shows that the point-estimate parameters learned in this way support Bayesian computation in representation space: value manifolds, key frames, and query trajectories implement Bayesian belief updates in context. The present paper explains why cross-entropy and gradient descent naturally create these structures.
@@ 6.1. Value Manifold Unfurling
In wind-tunnel experiments (Agarwal et al., 2025), we observe that:
@@ Results.
Figures Figure 7Figure 9 plot loss, accuracy, and entropy as a function of training step. As expected, EM and SGD methods converge to the same solution. This may not always happen due to local minima. EM converges much faster, with loss and entropy dropping sharply and accuracy rising quickly, while SGD progresses more slowly. This happens since $v_{j}$ are optimized directly in the function space, unlike SGD which optimizes in the parameter space. The argument is similar to functional gradient boost that converges faster than parameter updates (@(Friedman, 2001)). It is also interesting to note that even with this single layer and head architecture, the final empirical predictive entropy is very close to the theoretically minimum Bayesian entropy as described in Table 1. The theoretical entropy is higher due to addition of noise in defining $x_{t+1}$.
@@ 9.1. Bayesian Interpretations of Transformers
Several works argue that transformers implement approximate Bayesian inference, either behaviorally or via probing (e.g. Xie et al., 2022; von Oswald et al., 2023). Our companion paper (Agarwal et al., 2025) demonstrates exact Bayesian behavior and geometric signatures in small wind tunnels, while a scaling paper shows similar patterns in production LLMs. The present work explains how gradient dynamics produce these geometries. It further shows that minimum theoretical Bayesian predictive entropy is close to the empirical entropy. This is studied in greater detail in the companion paper.
@@ 9.2. Mechanistic Interpretability
Mechanistic interpretability studies identify specific heads and circuits performing copy, induction, and other algorithmic tasks (Elhage et al., 2021; Olsson et al., 2022). Our framework complements this line by explaining how specialization arises from the interaction between routing and content, rather than treating specialized heads as primitive.
@@ 9.3. Optimization and Implicit Bias
The implicit bias of gradient descent in linear and deep networks has been extensively studied (Soudry et al., 2018; Ba et al., 2022). We extend these ideas to attention: gradient descent implicitly favors representations where routing aligns with error geometry and values lie on low-dimensional manifolds that support Bayesian updating.